# GP-GPT: Large Language Model for Gene-Phenotype Mapping

### v3

@@ 1 Introduction
The relationship between genes, phenotypes and diseases is of fundamental importance but has not yet been fully understood. Due to the complex interplays and mutual regulations among genes, proteins, metabolomics, and phenotypes, finding the proper representations of their relationship has become a significant problem. There are numerous studies that focus on single-molecule or dual-molecule biological levels such as gene mutation, proteomic, gene expression, transcriptional regulation, pathway analysis, and clinical observation. These studies aim to address the questions in specific areas using simplified modelling. Existing research in human populations has elucidated the landscape of complex genomic relations. For instance, Genome-wide association study (GWAS) [1, 2, 3] is a well-established and effective approach for statistically discovering genetic loci associated with common diseases. Meta-analysis of genetic association provides new insights into cohorts of independent gene–disease associations study by combining them and resolving the discrepancies [4]. Genome annotation aims at understanding specific functional and structural properties which were conferred by nucleotide regions and amino acid sequences [5]. The Gene Ontology (GO) knowledge base (http://geneontology.org) comprehensively mapped the gene, functions of genes, and gene products into a formatted network [6] while human-specific gene networks like HumanNet v2 (http://www.inetbio.org/humannet) comprises a hierarchy of networks about gene-disease-drug associations [7]. The Online Mendelian Inheritance in Man(OMIM) [8] and DisGeNET [9] are essential knowledge platforms for disease genomics. Such large-scale genomics datasets and relevant studies have significantly expanded the scope of representative analysis of genes and diseases. However, holistic modeling at the whole genome-system scale remains challenging.
@@
To better address this fundamental problem of representing and mapping genes and related phenotypes/diseases, the ideal approach should consider the entire genome system collectively across various molecular and biological levels. Additionally, it should incorporate experimental knowledge, bio-computational evidence, and biological text data, which have been overlooked in previous studies despite their considerable value. Large language models(LLMs) have emerged as transformative tools in natural language processing, demonstrating unprecedented capabilities in understanding and generating human-like text [10, 11, 12, 12]. These models, trained on extensive corpora, have shown remarkable versatility across diverse domains [13, 14, 15, 16, 17, 18, 19]. There are also significant successes in applying language models to the general biomedical field [20, 21, 22, 23, 24, 25]. LLM’s potential to process and analyze complex, unstructured data presents a unique opportunity to address long-standing challenges in bioinformatics. The ability of LLMs to capture intricate relationships and context from text aligns well with the complexities inherent in genetic and phenotype data. This synergy opens new avenues for knowledge discovery, hypothesis generation, and the elucidation of previously obscure gene-disease associations.
@@
In this study, we focus on transforming the aforementioned problem into an AI-based system based on a gene-phenotype large language model, named GP-GPT. The main goal of GP-GPT is to integrate multiple sources of structured and unstructured genomic knowledge into a general LLM framework. Particularly, the model leverages both structured and unstructured data from multiple main data sources from OMIM, DisGeNET, and dbGaP [26]. By converting the data into bio-text, we identified and categorized gene entities, gene functions, protein entities, protein functions, phenotype entities, genotype-phenotype association analysis, and related biological mechanisms. These components were then integrated into an informative context, which was subsequently used to develop the GP-GPT in the form of generative language models. Therefore, the model considers multiple levels of bio-factors and achieves knowledge mapping of the entire genome system rather than individual relationships.
@@ 2.1 Language models for Knowledge Embedding and Relation Extraction
Recently, language models such as ChatGPT and GPT-4 [11], have shown remarkable capabilities in accelerating research innovation across natural sciences. ChatGPT, GPT-4, GPT-4V, and GPT-4 Turbo, developed by OpenAI, are prominent language models based on the autoregressive transformer decoder architecture, which utilizes self-attention mechanisms. These models, trained on extensive text corpora, excel in generating human-like text and demonstrate profound comprehension across diverse knowledge domains. In contrast, open-source models like Meta’s LLaMA [12] and Llama2 [12], which feature up to 70B parameters in their largest versions, aim to enhance training efficiency and reduce computational costs. Similarly, BLOOM (176B version) [27] and BLOOMZ (176B version) [28] emphasize linguistic versatility, highlighting a commitment to openness and accessibility. The more recent Falcon [29], with 40B parameters in its largest configuration, has also shown excellent performance on various benchmarks. These models exemplify the cutting-edge of universal language model research. LLaMA models, in particular, have performed exceptionally well on diverse benchmarks, making them some of the most popular open language models to date, hence their selection for analysis in this paper.
@@
Biomedical natural language processing enables the automatic extraction of crucial information from medical literature, including insights into genetic diseases and associated variants. A key task in this field is Biological Relation Extraction (RE), which involves identifying relationships between two or more entities within specific contexts, which is crucial to identify biological entities such as genes, proteins, diseases, drugs, and miRNAs in textual data [30, 9, 31, 32]. Biological RE models are widely used to extract knowledge from various sources, including medical literature. A key task in Biological RE is identifying relationships between two or more entities within specific contexts. To advance NLP and machine learning techniques for biomedical RE, considerable effort has been invested in developing relevant corpora [33, 34, 35, 36]. Unlike simple regular expression-based methods, Biological RE models leverage the entire context of a sentence to accurately identify and classify these entities, overcoming the limitations of more traditional approaches. Although machine learning (ML) techniques generally outperform traditional methods, they require substantial amounts of manually annotated training data. This presents a significant challenge, especially in the genomics domain, where labelling training sequences at the individual word level is time-consuming and requires domain-specific expertise. Owing to large language models which offers a solution by enabling fine-tuning for Biological RE tasks with minimal training data, thereby reducing the burden of extensive manual annotation.
@@ 2.2 Instruction Fine-tuning and Parameter efficient Training
Instruction fine-tuning has emerged as a pivotal technique in improving the performance and generalization of language models across various tasks. The work by Wei et al. [37] introduces a novel training methodology that leverages instruction fine-tuning to enhance model capabilities. By fine-tuning models on multiple tasks with carefully curated instructions, this approach not only improves task-specific performance but also enhances the model’s ability to generalize across different domains.
The Stanford CRFM report [38] further substantiates the efficacy of instruction fine-tuning. The report provides a comprehensive evaluation of different models and training strategies, demonstrating that instruction fine-tuning can achieve significant improvements in performance even with limited computational resources. This underscores the potential of instruction fine-tuning as a cost-effective solution for developing versatile language models.
Parameter-efficient training methods aim to optimize model performance while minimizing computational overhead [39]. Recent advancements in this field are well-documented in the work by Liu et al. [38], which explores various parameter-efficient techniques [40]. These techniques are designed to reduce the number of trainable parameters without compromising model accuracy, making them particularly suitable for environments with constrained resources .
Key innovations include low-rank adaptation, pruning, and quantization, which collectively contribute to maintaining model efficacy with fewer parameters. Experimental results in Liu et al. [41] highlight that such methods can achieve near state-of-the-art performance with a fraction of the computational cost, thereby promoting sustainable AI development practices.
Combining instruction fine-tuning with parameter-efficient training presents a promising avenue for future research [42]. Integrating these methods can potentially lead to the development of highly adaptable and resource-efficient models. The synergy between these approaches lies in their complementary strengths: instruction fine-tuning enhances task adaptability, while parameter-efficient training ensures computational feasibility.
@@ 2.3 Language Model Applications in Bioinformatics
In recent years, the emergence of language models for bioinformatics has shown significant progress, with notable examples including AlphaFold2 [43], AlphaFold3  [44], GeneGPT [45], BioT5 [46], BioT5+ [47], and Med-Gemini [48]. These models can be categorized into single-domain and multi-domain approaches, each leveraging different strategies for pretraining and application in the field.
@@
Single-domain approaches pretrain their models solely on tokens from a specific domain. A prominent example is GeneGPT [45], which introduces a novel method for teaching large language models (LLMs) to utilize the Web APIs of the National Center for Biotechnology Information (NCBI) to answer genomics-related questions. By focusing exclusively on genomics data, GeneGPT can provide precise and relevant responses tailored to this domain.
@@
Based on different foundational models, large language models in the bio-informatics domain can be categorized into BERT-based LLMs, T5, GPT and others [49]. BERT-based large models related to biology include the SciBERT model. SciBERT is pre-trained on scientific publications from Semantic Scholar, mainly using the WordPiece unsupervised tokenization technique [50]. ClinicalBERT is a pre-trained model built on a BERT-based architecture using clinical notes as training data. The types of clinical notes include nursing, radiology, and others [51]. SciFive is a T5-based generative model pre-trained on millions of PubMed abstracts and PMC scientific publications. This model excels at handling biomedical tasks such as question answering and named entity recognition, outperforming models like BioBERT, T5, and BERT [52]. BioGPT is a GPT-2-based model pre-trained on 15 million PubMed articles, making it one of the first GPT-inspired biomedical frameworks. At the time, it was also one of the top-performing models in its domain [53].
@@
In the application of large language models to biological sequence data, research can be broadly categorized into four areas: DNA sequences, RNA sequences, protein sequences, and multi-omics sequencing data. Each of these areas presents unique challenges and opportunities for decoding and understanding biological systems. For DNA sequences, models such as Enformer [54], Nucleotide Transformer [55], GenSLMs [56], and HyenaDNA [57] have been developed to decipher the genetic blueprint of life, advancing our understanding of gene regulation, function, and expression. Similarly, in the study of ribonucleic acid (RNA), foundational models like RNABERT [58], RNA-FM [59] are instrumental in unravelling the complex mechanisms that drive RNA-mediated biological processes. These models play a crucial role in areas such as transcriptomics and gene expression analysis. In the protein domain, where the focus is on protein structure, function, and interactions, numerous models have been proposed, such as AlphaFold2 [43]. The model aims to predict protein folding, dynamics, and interactions, all of which are essential for understanding cellular processes and developing therapeutic strategies. Multi-omics sequencing data integrates various biological layers and enables the analysis of more complex and comprehensive downstream tasks. Models like Geneformer  [60], scFoundation  [61], AlphaFold3 [44], and scGPT  [62] is at the forefront of this domain, offering powerful tools for integrating single-cell data and other omics modalities to gain a holistic view of biological systems.
@@
On the other hand, multi-domain approaches employ a diverse array of data modalities during pretraining. BioT5+, for instance, utilizes molecular structures, textual descriptions, and molecular naming conventions [47]. This approach enhances the model’s ability to perform a wider range of downstream tasks. Med-Gemini extends this concept further by incorporating an even broader spectrum of multimodal data, including 2D and 3D radiology images, histopathology patches, ophthalmology images, dermatology images, and genetic data [48]. This extensive pretraining allows Med-Gemini to address a vast array of biomedical tasks.
@@
Furthermore, these models often function through an interactive interface, providing chatbot-like functionalities for downstream tasks. Models like BioT5 [46] enable users to engage in intuitive dialogue formats, posing inquiries in plain text and receiving detailed, contextually relevant responses. This conversational ability allows chemical LLMs to perform complex tasks that require understanding and reasoning beyond simple query responses. The majority of these models achieve their functionality through two primary methods: fine-tuning on domain-specific data or directly interfacing with existing APIs to provide responses. For example, GeneGPT [45] uses carefully designed prompts to demonstrate to the LLM how to query APIs effectively, utilizing appropriate APIs during the inference phase to respond to user requests. Alternatively, models like Med-Gemini can directly answer user queries through the model’s outputs, demonstrating versatility and a comprehensive understanding of a wide range of topics [48].
@@ 2.4 Mapping of Genotypes and Phenotypes
Genotype-to-phenotype mapping is a traditional work in bioinformatics. In a statistical view, GWAS [1, 2, 3] have played a dominant role in discovering potential relations between genotypes and phenotypes on a population scale. It has tested hundreds of thousands of genetic variants across human and other genomes to find those statistically associated traits or diseases, through which gaining insight into a phenotype’s underlying biology, estimating its heritability, and discovering genetic correlations. From a micro view, advancements in sequencing technologies, particularly single-cell RNA-seq (scRNA-seq), have significantly advanced our understanding. These techniques expose the intricate dynamics of gene expression at the cellular level, unveiling the extensive landscape of genotypes shaped by numerous factors [63]. In recent years, the utilization of machine learning, particularly the adoption of self-supervised transformer models originally developed for natural language processing (NLP), has demonstrated significant potential in the analysis of complex biological datasets  [60, 62].
At the cohort level, meta-analysis is an effective approach to integrate multiple association studies of genes and phenotypes. Meta-analysis is a statistical tool used to combine results from different studies on the same topic and quantitative synthesis of research findings. Meta-analysis enables the identification of genuine associations between genes and phenotype [64]. Particularly for the studies based on human populations, a comprehensive catalogue of all published GWAS and association results have become publicly available, which provided resources for human-specific meta-analysis [4].
At the level of human cells, there is a principal goal of revealing the landscape of a combination of genes that results in an extraordinarily diverse phenomenon. Phenomic databases [65], have been collected as comprehensively as systematically assemblages of phenotype information.Freimer, et. al. [66] has presented a computational integrated genetics framework designed to analyze and interpret the high-dimensional landscape of genotypes and their associated phenotype. The study introduces the use of self-supervised language models to simultaneously map the genotype-phenotype landscape, proposing the concept of integrated genetics as an alternative to traditional forward and reverse genetics. In this approach, phenotypes such as sex, age, anatomical tissue, and cell types are integrated and learned alongside genotypic data obtained from scRNA-seq. This method deepens our understanding of the biological context of gene expression and, ultimately, the genotype-phenotype relationship.

### v4

@@ 1 Introduction
The relationship between genes, phenotypes and diseases is of fundamental importance but has not yet been fully understood. Due to the complex interplays and mutual regulations among genes, proteins, metabolomics, and phenotypes, finding the proper representations of their relationship has become a significant problem. There are numerous studies that focus on single-molecule or dual-molecule biological levels such as gene mutation, proteomic, gene expression, transcriptional regulation, pathway analysis, and clinical observation. These studies aim to address the questions in specific areas using simplified modelling. Existing research in human populations has elucidated the landscape of complex genomic relations. For instance, Genome-wide association study (GWAS) [visscher201710, watanabe2019global, jansen2019genome] is a well-established and effective approach for statistically discovering genetic loci associated with common diseases. Meta-analysis of genetic association provides new insights into cohorts of independent gene–disease associations study by combining them and resolving the discrepancies [macarthur2017new]. Genome annotation aims at understanding specific functional and structural properties which were conferred by nucleotide regions and amino acid sequences [abril2019genome]. The Gene Ontology (GO) knowledge base (http://geneontology.org) comprehensively mapped the gene, functions of genes, and gene products into a formatted network [gene2023gene] while human-specific gene networks like HumanNet v2 (http://www.inetbio.org/humannet) comprises a hierarchy of networks about gene-disease-drug associations [hwang2019humannet]. The Online Mendelian Inheritance in Man(OMIM) [amberger2015omim] and DisGeNET [pinero2020disgenet] are essential knowledge platforms for disease genomics. Such large-scale genomics datasets and relevant studies have significantly expanded the scope of representative analysis of genes and diseases. However, holistic modeling at the whole genome-system scale remains challenging.
@@
To better address this fundamental problem of representing and mapping genes and related phenotypes/diseases, the ideal approach should consider the entire genome system collectively across various molecular and biological levels. Additionally, it should incorporate experimental knowledge, bio-computational evidence, and biological text data, which have been overlooked in previous studies despite their considerable value. Large language models(LLMs) have emerged as transformative tools in natural language processing, demonstrating unprecedented capabilities in understanding and generating human-like text [raffel2020exploring, achiam2023gpt, touvron2023llama, touvron2023llama]. These models, trained on extensive corpora, have shown remarkable versatility across diverse domains [liu2023summary, zhao2023brain, ma2024iterative, dai2023chataug, liu2023radiology, liao2023differentiating, liu2023context]. There are also significant successes in applying language models to the general biomedical field [zhang2023biomedgpt, dai2023ad, holmes2023evaluating, rezayi2022clinicalradiobert, wang2023prompt, liu2023holistic]. LLM’s potential to process and analyze complex, unstructured data presents a unique opportunity to address long-standing challenges in bioinformatics. The ability of LLMs to capture intricate relationships and context from text aligns well with the complexities inherent in genetic and phenotype data. This synergy opens new avenues for knowledge discovery, hypothesis generation, and the elucidation of previously obscure gene-disease associations.
@@
In this study, we focus on transforming the aforementioned problem into an AI-based system based on a gene-phenotype large language model, named GP-GPT. The main goal of GP-GPT is to integrate multiple sources of structured and unstructured genomic knowledge into a general LLM framework. Particularly, the model leverages both structured and unstructured data from multiple main data sources from OMIM, DisGeNET, and dbGaP [mailman2007ncbi]. By converting the data into bio-text, we identified and categorized gene entities, gene functions, protein entities, protein functions, phenotype entities, genotype-phenotype association analysis, and related biological mechanisms. These components were then integrated into an informative context, which was subsequently used to develop the GP-GPT in the form of generative language models. Therefore, the model considers multiple levels of bio-factors and achieves knowledge mapping of the entire genome system rather than individual relationships.
@@ 2.1 Language models for Knowledge Embedding and Relation Extraction
Recently, language models such as ChatGPT and GPT-4 [achiam2023gpt], have shown remarkable capabilities in accelerating research innovation across natural sciences. ChatGPT, GPT-4, GPT-4V, and GPT-4 Turbo, developed by OpenAI, are prominent language models based on the autoregressive transformer decoder architecture, which utilizes self-attention mechanisms. These models, trained on extensive text corpora, excel in generating human-like text and demonstrate profound comprehension across diverse knowledge domains. In contrast, open-source models like Meta’s LLaMA [touvron2023llama] and Llama2 [touvron2023llama], which feature up to 70B parameters in their largest versions, aim to enhance training efficiency and reduce computational costs. Similarly, BLOOM (176B version) [le2023bloom] and BLOOMZ (176B version) [muennighoff2022crosslingual] emphasize linguistic versatility, highlighting a commitment to openness and accessibility. The more recent Falcon [almazrouei2023falcon], with 40B parameters in its largest configuration, has also shown excellent performance on various benchmarks. These models exemplify the cutting-edge of universal language model research. LLaMA models, in particular, have performed exceptionally well on diverse benchmarks, making them some of the most popular open language models to date, hence their selection for analysis in this paper.
@@
Biomedical natural language processing enables the automatic extraction of crucial information from medical literature, including insights into genetic diseases and associated variants. A key task in this field is Biological Relation Extraction (RE), which involves identifying relationships between two or more entities within specific contexts, which is crucial to identify biological entities such as genes, proteins, diseases, drugs, and miRNAs in textual data [gill2024large, pinero2020disgenet, babbi2017edgar, marchesin2022tbga]. Biological RE models are widely used to extract knowledge from various sources, including medical literature. A key task in Biological RE is identifying relationships between two or more entities within specific contexts. To advance NLP and machine learning techniques for biomedical RE, considerable effort has been invested in developing relevant corpora [kim2003genia, pyysalo2007bioinfer, krallinger2008overview, luo2022biored]. Unlike simple regular expression-based methods, Biological RE models leverage the entire context of a sentence to accurately identify and classify these entities, overcoming the limitations of more traditional approaches. Although machine learning (ML) techniques generally outperform traditional methods, they require substantial amounts of manually annotated training data. This presents a significant challenge, especially in the genomics domain, where labelling training sequences at the individual word level is time-consuming and requires domain-specific expertise. Owing to large language models which offers a solution by enabling fine-tuning for Biological RE tasks with minimal training data, thereby reducing the burden of extensive manual annotation.
@@ 2.2 Instruction Fine-tuning and Parameter efficient Training
Instruction fine-tuning has emerged as a pivotal technique in improving the performance and generalization of language models across various tasks. The work by Wei et al. [wu2023interpretability] introduces a novel training methodology that leverages instruction fine-tuning to enhance model capabilities. By fine-tuning models on multiple tasks with carefully curated instructions, this approach not only improves task-specific performance but also enhances the model’s ability to generalize across different domains.
The Stanford CRFM report [stanfordStanfordCRFM] further substantiates the efficacy of instruction fine-tuning. The report provides a comprehensive evaluation of different models and training strategies, demonstrating that instruction fine-tuning can achieve significant improvements in performance even with limited computational resources. This underscores the potential of instruction fine-tuning as a cost-effective solution for developing versatile language models.
Parameter-efficient training methods aim to optimize model performance while minimizing computational overhead [hu2021lora]. Recent advancements in this field are well-documented in the work by Liu et al. [stanfordStanfordCRFM], which explores various parameter-efficient techniques [houlsby2019parameter]. These techniques are designed to reduce the number of trainable parameters without compromising model accuracy, making them particularly suitable for environments with constrained resources .
Key innovations include low-rank adaptation, pruning, and quantization, which collectively contribute to maintaining model efficacy with fewer parameters. Experimental results in Liu et al. [ouyang2022training] highlight that such methods can achieve near state-of-the-art performance with a fraction of the computational cost, thereby promoting sustainable AI development practices.
Combining instruction fine-tuning with parameter-efficient training presents a promising avenue for future research [dettmers2024qlora]. Integrating these methods can potentially lead to the development of highly adaptable and resource-efficient models. The synergy between these approaches lies in their complementary strengths: instruction fine-tuning enhances task adaptability, while parameter-efficient training ensures computational feasibility.
@@ 2.3 Language Model Applications in Bioinformatics
In recent years, the emergence of language models for bioinformatics has shown significant progress, with notable examples including AlphaFold2 [bryant2022improved], AlphaFold3 [abramson2024accurate], GeneGPT [jin2024genegpt], BioT5 [pei2023biot5], BioT5+ [pei2024biot5+], and Med-Gemini [yang2024advancing]. These models can be categorized into single-domain and multi-domain approaches, each leveraging different strategies for pretraining and application in the field.
@@
Single-domain approaches pretrain their models solely on tokens from a specific domain. A prominent example is GeneGPT [jin2024genegpt], which introduces a novel method for teaching large language models (LLMs) to utilize the Web APIs of the National Center for Biotechnology Information (NCBI) to answer genomics-related questions. By focusing exclusively on genomics data, GeneGPT can provide precise and relevant responses tailored to this domain.
@@
Based on different foundational models, large language models in the bio-informatics domain can be categorized into BERT-based LLMs, T5, GPT and others [bi2024ai]. BERT-based large models related to biology include the SciBERT model. SciBERT is pre-trained on scientific publications from Semantic Scholar, mainly using the WordPiece unsupervised tokenization technique [beltagy2019scibert]. ClinicalBERT is a pre-trained model built on a BERT-based architecture using clinical notes as training data. The types of clinical notes include nursing, radiology, and others [alsentzer2019publicly]. SciFive is a T5-based generative model pre-trained on millions of PubMed abstracts and PMC scientific publications. This model excels at handling biomedical tasks such as question answering and named entity recognition, outperforming models like BioBERT, T5, and BERT [phan2021scifive]. BioGPT is a GPT-2-based model pre-trained on 15 million PubMed articles, making it one of the first GPT-inspired biomedical frameworks. At the time, it was also one of the top-performing models in its domain [luo2022biogpt].
@@
In the application of large language models to biological sequence data, research can be broadly categorized into four areas: DNA sequences, RNA sequences, protein sequences, and multi-omics sequencing data. Each of these areas presents unique challenges and opportunities for decoding and understanding biological systems. For DNA sequences, models such as Enformer [avsec2021effective], Nucleotide Transformer [dalla2023nucleotide], GenSLMs [zvyagin2023genslms], and HyenaDNA [nguyen2306hyenadna] have been developed to decipher the genetic blueprint of life, advancing our understanding of gene regulation, function, and expression. Similarly, in the study of ribonucleic acid (RNA), foundational models like RNABERT [akiyama2022informative], RNA-FM [chen2022interpretable] are instrumental in unravelling the complex mechanisms that drive RNA-mediated biological processes. These models play a crucial role in areas such as transcriptomics and gene expression analysis. In the protein domain, where the focus is on protein structure, function, and interactions, numerous models have been proposed, such as AlphaFold2 [bryant2022improved]. The model aims to predict protein folding, dynamics, and interactions, all of which are essential for understanding cellular processes and developing therapeutic strategies. Multi-omics sequencing data integrates various biological layers and enables the analysis of more complex and comprehensive downstream tasks. Models like Geneformer [theodoris2023transfer], scFoundation [hao2024large], AlphaFold3 [abramson2024accurate], and scGPT [cui2024scgpt] is at the forefront of this domain, offering powerful tools for integrating single-cell data and other omics modalities to gain a holistic view of biological systems.
@@
On the other hand, multi-domain approaches employ a diverse array of data modalities during pretraining. BioT5+, for instance, utilizes molecular structures, textual descriptions, and molecular naming conventions [pei2024biot5+]. This approach enhances the model’s ability to perform a wider range of downstream tasks. Med-Gemini extends this concept further by incorporating an even broader spectrum of multimodal data, including 2D and 3D radiology images, histopathology patches, ophthalmology images, dermatology images, and genetic data [yang2024advancing]. This extensive pretraining allows Med-Gemini to address a vast array of biomedical tasks.
@@
Furthermore, these models often function through an interactive interface, providing chatbot-like functionalities for downstream tasks. Models like BioT5 [pei2023biot5] enable users to engage in intuitive dialogue formats, posing inquiries in plain text and receiving detailed, contextually relevant responses. This conversational ability allows chemical LLMs to perform complex tasks that require understanding and reasoning beyond simple query responses. The majority of these models achieve their functionality through two primary methods: fine-tuning on domain-specific data or directly interfacing with existing APIs to provide responses. For example, GeneGPT [jin2024genegpt] uses carefully designed prompts to demonstrate to the LLM how to query APIs effectively, utilizing appropriate APIs during the inference phase to respond to user requests. Alternatively, models like Med-Gemini can directly answer user queries through the model’s outputs, demonstrating versatility and a comprehensive understanding of a wide range of topics [yang2024advancing].
@@ 2.4 Mapping of Genotypes and Phenotypes
Genotype-to-phenotype mapping is a traditional work in bioinformatics. In a statistical view, GWAS [visscher201710, watanabe2019global, jansen2019genome] have played a dominant role in discovering potential relations between genotypes and phenotypes on a population scale. It has tested hundreds of thousands of genetic variants across human and other genomes to find those statistically associated traits or diseases, through which gaining insight into a phenotype’s underlying biology, estimating its heritability, and discovering genetic correlations. From a micro view, advancements in sequencing technologies, particularly single-cell RNA-seq (scRNA-seq), have significantly advanced our understanding. These techniques expose the intricate dynamics of gene expression at the cellular level, unveiling the extensive landscape of genotypes shaped by numerous factors [norman2019exploring]. In recent years, the utilization of machine learning, particularly the adoption of self-supervised transformer models originally developed for natural language processing (NLP), has demonstrated significant potential in the analysis of complex biological datasets [theodoris2023transfer, cui2024scgpt].
At the cohort level, meta-analysis is an effective approach to integrate multiple association studies of genes and phenotypes. Meta-analysis is a statistical tool used to combine results from different studies on the same topic and quantitative synthesis of research findings. Meta-analysis enables the identification of genuine associations between genes and phenotype [trikalinos2008meta]. Particularly for the studies based on human populations, a comprehensive catalogue of all published GWAS and association results have become publicly available, which provided resources for human-specific meta-analysis [macarthur2017new].
At the level of human cells, there is a principal goal of revealing the landscape of a combination of genes that results in an extraordinarily diverse phenomenon. Phenomic databases [freimer2003human], have been collected as comprehensively as systematically assemblages of phenotype information.Freimer, et. al. [khodaee2024multimodal] has presented a computational integrated genetics framework designed to analyze and interpret the high-dimensional landscape of genotypes and their associated phenotype. The study introduces the use of self-supervised language models to simultaneously map the genotype-phenotype landscape, proposing the concept of integrated genetics as an alternative to traditional forward and reverse genetics. In this approach, phenotypes such as sex, age, anatomical tissue, and cell types are integrated and learned alongside genotypic data obtained from scRNA-seq. This method deepens our understanding of the biological context of gene expression and, ultimately, the genotype-phenotype relationship.